# Domain Generalizable Knowledge Tracing via Concept Aggregation and Relation-Based Attention

### v1

@@ Abstract
Knowledge Tracing (KT) is a critical task in online education systems, aiming to monitor students’ knowledge states throughout a learning period. Common KT approaches involve predicting the probability of a student correctly answering the next question based on their exercise history. However, these methods often suffer from performance degradation when faced with the scarcity of student interactions in new education systems. To address this, we leverage student interactions from existing education systems to mitigate performance degradation caused by limited training data. Nevertheless, these interactions exhibit significant differences since they are derived from different education systems. To address this issue, we propose a domain generalization approach for knowledge tracing, where existing education systems are considered source domains, and new education systems with limited data are considered target domains. Additionally, we design a domain-generalizable knowledge tracing framework (DGKT) that can be applied to any KT model. Specifically, we present a concept aggregation approach designed to reduce conceptual disparities within sequences of student interactions from diverse domains. To further mitigate domain discrepancies, we introduce a novel normalization module called Sequence Instance Normalization (SeqIN). Moreover, to fully leverage exercise information, we propose a new knowledge tracing model tailored for the domain generalization KT task, named Domain-Generalizable Relation-based Knowledge Tracing (DGRKT). Extensive experiments across five benchmark datasets demonstrate that the proposed method performs well despite limited training data. Traditional KT models show significant performance degradation on target domains, while models with the DGKT framework alleviate this issue. Our proposed DGRKT model outperforms five knowledge tracing models, achieving an average AUC improvement of 4.16%, benefiting from a specially designed relation-based attention mechanism.
Index Terms:
@@
Knowledge tracing, intelligent education, domain generalization, knowledge concept.
I Introduction
@@
Over the past decades, a large number of online education systems have emerged, which provide remote learning environments and personalized guidance for users. These advancements have revolutionized the educational landscape, enabling platforms to cater to the individual learning needs of students more effectively [1, 2, 3]. Knowledge tracing (KT) plays a crucial role in this context, as it allows education systems to monitor and evaluate the evolving knowledge states of students. Knowledge tracing (KT) aims to trace a student’s knowledge states accurately which is often realized via the task of predicting students’ future performance according to their historical interactions [4, 5]. This enables online education platforms to better assess students’ comprehension and provide tailored assistance for learners [6, 7].
A number of KT models have demonstrated their effectiveness  [8, 9, 10]. However, these methods require a substantial amount of student interactions to adequately train a KT model for a specific subject. In the majority of knowledge tracing datasets, students’ problem-solving records typically have a scale of several hundred thousand or more. For example, there are 346, 860 interactions in ASSISTment 2009 which is completely enough to train a KT model sufficiently while it is difficult for a newly developed education system to gain such a substantial amount of data in the beginning. In practice, developing a new online education system or adding a new question bank to an existing one often lacks the availability of abundant student interactions, leading to a scarcity of data for training a KT model [11, 12]. This scarcity of interactions results in the challenge of insufficient training data. As shown in Fig. 1, common KT models suffer significant AUC degradation when confronted with limited data sizes.
Figure 1: The AUC result for diffarent KT models with different datasize.
@@
Figure 2: An intuitive explanation for DGKT. Existing systems contain abundant interaction records while new education system only a few interactions. We aim to assess student’s knowledge state and predict the answer correctness despite lack of the interaction data.
Currently, few studies have delved into the data scarcity issue in knowledge tracing. Some prior efforts have aimed to address the sparse problem that students tend to interact with only a small set of questions, using approaches like pre-trained question embeddings [13] or contrastive learning [7, 14, 15]. However, these methods assume relatively sufficient interaction records. In cases where interactions are severely limited, these approaches might not be effective in learning question embeddings and students’ knowledge states, as they heavily rely on a larger amount of student interactions. Thus, the scarcity of student interactions in new educational systems presents a significant challenge.
@@
In the absence of student interaction records in the new system, numerous student interaction records from diverse online education platforms offer valuable insights. These records may originate from different sources, yet they mirror the process of knowledge acquisition. Extracting overarching cognitive patterns from these varied student interaction sequences is pivotal for addressing the data scarcity issue in KT tasks. AdaptKT, as proposed by [16], delved into domain adaptation for knowledge tracing. However, it is demanding that AdaptKT requires the question texts and relatively sufficient interactions in the target domain, which is often not met in reality. Moreover interaction sequences from different education system differs a lot so it is unfeasible to directly use them as training data. Thus, the significant differences between various educational systems present another major challenge.
We innovatively frame the dilemma of the new education system as a domain generalization issue, as shown in Fig. 2. We aim to design a novel KT method that derives meta-knowledge from various student interaction sequences from different education systems(source domains), enabling seamless transfer to a new education system(target domain). In the context of domain generalization for knowledge tracing, different domains corresponds to the questions and interactions from different online education platforms, such as engineering statics course and Algebra course. These domains have different knowledge concepts and these concepts are related to different questions, we aim to assess student’s mastery on various concepts. Sufficient student interaction records from specific systems are treated as source domains, while a limited number of interactions in the new system represent the target domain.
@@
To solve the domain generalization issue, this paper introduces a novel approach, namely Domain Generalizable Knowledge Tracing (DGKT). DGKT’s goal is twofold: firstly, to utilize data from multiple source domains for training a versatile KT model, and subsequently, to swiftly adapt this model to a novel target domain with a few student interactions. Notably, DGKT faces two primary challenges: 1) significant differences of concepts among the source domains, and 2) scarce student interactions within the target domain.
For the problem of significant differences of concepts among the source domains, since diverse source domains encompass different concepts, raw concepts are ineffective to achieve domain generalization. To overcome this challenge, we propose the concept aggregation—a clustering algorithm for concept embedding that analyzes students’ interaction sequences from a unified perspective. Moreover, we propose a relation-based attention encoder that fully leverage the relation
@@
For the problem of scarce student interactions within the target domain, we find it difficult to learn the target question embedding with a few interactions. Consequently, we design a unique concept representation for target domain that can effectively adapt to target domain with a few training data. Moreover, we propose a relation-based attention that fully leverage exercise information in target domain. The contributions of this paper are summarized as follows:
  • A novel solution of knowledge tracing for the data scarcity issue. The data scarcity issue is valuable but insufficiently studied. We aim to train a domain-generalizable KT model from auxiliary source domains and quickly adapt the model to new target domains with only a few interactions.
@@
  • Innovative concept aggregation for reducing domain discrepancy. This addresses the challenge that questions and concepts from different domains have enormous differences. Through concept aggregation, similar concepts are associated with the same cluster centroid.
@@
  • Domain-generalizable relation-based knowledge tracing. We introduce a domain-generalizable relation-based knowledge tracing approach that leverages relational information within the target domain. Our specially designed model demonstrates superior performance in domain generalization for knowledge tracing.
@@
After conducting extensive experiments on five benchmark datasets, our approach shows a performance of 4.16% improvement over five KT methods on domain generalization of knowledge tracing tasks.
II Related Work
In this section, we introduce the related work including congnitive diagnosis, knwoledge tracing and domain generalization.
II-A Cognitive Diagnosis
In online education systems, cognitive diagnosis is an essential task aimed at analyzing a student’s performance and providing a detailed evaluation of their proficiency across various concepts. Over the past decades, numerous cognitive diagnosis models have been developed, which can be broadly classified into two categories: traditional cognitive diagnosis models and deep learning-based cognitive diagnosis models. Traditional models include IRT [17], DINA [18] and MIRT [19], and one of the most prominent is IRT. IRT represents a student’s ability as a one-dimensional scalar and calculates the probability of correctly answering an exercise using a logistic function. This model has been widely used due to its simplicity and interpretability. However, it is limited by its assumption of unidimensionality and its relatively simplistic modeling of student responses. In contrast, deep learning-based cognitive diagnosis models leverage the power of neural networks to capture more complex patterns in student data. A notable example is NeuralCD [20], which applies a multilayer perceptron to model the cognitive diagnosis process. By utilizing deep neural networks, NeuralCD can learn intricate relationships between student performance and underlying knowledge concepts.
While cognitive diagnosis models are effective in evaluating a student’s proficiency at a given point in time, they often lack the capability to continuously track the student’s knowledge state over an extended learning period. This limitation hinders their ability to provide ongoing, dynamic feedback that adapts to the student’s evolving learning needs.
II-B Knowledge Tracing
@@
Recently, many deep models for KT have emerged. The earliest deep model for KT is Deep Knowledge Tracing (DKT) [6]. DKT applies recurrent neural networks (RNNs) and outperforms traditional KT models. Many variants of DKT are proposed afterwards such as DKT+ [24]. Exercise-Enhanced Recurrent Neural Network (EERNN) [25] is proposed for student performance prediction by taking full advantage of both student exercising records and the text of each exercise. Another typical deep KT model is Separated Self-AttentIve Neural Knowledge Tracing (SAINT) [8] which applies Transformer to KT tasks. Attentive Knowledge Tracing (AKT) [9] is another attention-based KT model using a novel monotonic attention mechanism that relates a learner’s future responses to assessment questions to their past responses. Learning Process-consistent Knowledge Tracing (LPKT) [26] monitors students’ knowledge state through directly modeling their learning process. As far as we know, few works focus on the data scarcity issue in online education systems, and we innovatively attempt to train a generalizable KT model.
II-C Domain Generalization
Domain generalization (DG) has attracted increasing interest in recent years [27]. Some DG methods improve generalization via designing novel learning strategies or augmentation [28, 29]. Mancini et al. use learnable weights for aggregating the predictions from different source-specific classifiers [30], where a domain predictor is adopted to predict the probability of a sample belonging to each domai. Learning and Removing Domain-specific features for Generalization (LRDG) [31] learns a domain-invariant model by tactically removing domain-specific features from the input images. There are several DG methods based on meta-learning. Inspired by model-agnostic meta-learning (MAML) [32], Li et al. propose Meta-Learning for DG (MLDG) [33], which is an optimization-based meat-learning strategy for domain generalization. MetaReg [34] introduce a meta regularizer in meta-learning for fine-grained regularization. Another way to solve the DG problem is representation learning. Many DG methods reduce distribution discrepancy across training domains to learn domain invariant representations via feature normalization [35] or minimize feature distribution divergence explicitly  [36]. SNR  [37] is proposed to simultaneously ensure both high generalization and discrimination capability of the networks. However, these methods cannot be directly borrowed for KT tasks since KT relies on comprehending a student’s learning trajectory.
III Domain-Generalizable Knowledge Tracing
@@
In this section, we elaborate on the proposed DGKT method. Firstly, we formulate the knowledge tracing (KT) tasks and their domain generalization setting. Subsequently, we present the model architecture and concept aggregation. Furthermore, we introduce the generalization of KT model, which empowers the model to acquire meta-knowledge from source domains and effectively apply it to the target domain.
III-A Problem Definition
@@
A KT task with few sequences. In a KT task, there are very few sequences of interactions for training, denoted as $I=\{(q_{1},r_{1}),(q_{2},r_{2}),...,(q_{T},r_{T})\}$. Here, $T$ represents the length of the sequence, $q_{t}\in\mathbb{N}^{+}$ corresponds to the question ID of the $t$-th interaction ($t\leq T$), and $r_{t}\in\{0,1\}$ means the correctness of the student’s answer to the question $q_{t}$. These questions are associated with $n_{c}$ concepts that students need to master. The objective is to train a KT model capable of mining the knowledge state of students and predicting the probability that a student will answer the next question correctly, denoted as $P(r_{n+1}|q_{n+1},I)$. Since data scarcity affects model training, we resort to other available sequences as auxiliary source domain data for domain generalization.
@@
Domain generalization setting. In the context of domain generalization for knowledge tracing, there are $N$ auxiliary source domains $\{D_{s}^{i}|1\leq i\leq N\}$ and a target domain $D_{t}$. Each source domain $D_{s}^{i}$ consists of $m_{i}$ sequences of student interactions, denoted as $D_{s}^{i}=\{I_{j}^{i}|1\leq j\leq m_{i}\}$ along with a set of questions $\{q_{1},q_{2},...,q_{n_{q_{i}}}\}$ and a set of concepts $\{c_{1},c_{2},...,c_{n_{c_{i}}}\}$, where $n_{i}$ represents the number of questions, $k_{i}$ represents the number of concepts. The target domain, on the other hand, contains only $m_{t}$ student interaction sequences, represented as $D_{t}=\{I_{j}^{t}|1\leq j\leq m_{t}\}$, with $m_{t}\ll m_{i}$. Importantly, the data $D_{t}$ from the target domain is unseen during the meta-training phase. The objective is to train a generalized KT model by leveraging all the source domain data $\{D_{s}^{i}|1\leq i\leq N\}$, and subsequently adapt this model to the target domain data $D_{t}$ for knowledge tracing.
@@
III-B Model Architecture
The process outlined in Fig. 3 illustrates the key steps of our DGKT approach. Initially, we employ the feature embedding module to transform the sequence into embedding sequence. These embedding sequences are then input into the knowledge state encoder, which generates the hidden knowledge state. Ultimately, this knowledge state is decoded by the knowledge state decoder, producing the anticipated probability of a student providing a correct response to the subsequent question.
Feature embedding module. We aim to convert the sequence of questions and responses into the embedding sequence. For each knowledge tracing domain, we have a concept matrix $\textbf{Q}\in\mathbb{R}^{n_{c}\times n_{q}}$, where $n_{q}$ represents the total number of question IDs and $n_{c}$ represents the total number of knowledge concept IDs. The element $\textbf{Q}_{ij}$ is set to 1 if the $j$-th question is related to the $i$-th knowledge concept. Thus, we learn the embedding $e_{q_{t}}\in\mathbb{R}^{d}$ of question $q_{t}$ ($d$ is the dimension of the embedding), which can be written by
$e_{q_{t}}=\frac{\sum_{i}\mathbb{I}(Q_{iq_{t}}=1)e_{i}}{\sum_{i}\mathbb{I}(Q_{iq_{t}}=1)},$ (1)
where $\mathbb{I}(\cdot)$ represents the indicator function, which equals 1 if the condition inside the parentheses is true, and 0 otherwise. $e_{i}\in\mathbb{R}^{d}$ is a learnable vector representing the i-th concept.
Using the questions and responses, we construct the question-response embedding $e_{qr_{t}}\in\mathbb{R}^{2d}$ as follows:

### v2

# Domain Generalizable Knowledge Tracing via Concept Aggregation and Relation-Aware Attention

@@ Abstract
Knowledge Tracing (KT) is a fundamental task in online education systems, aiming to model students’ knowledge states by predicting their future performance. However, in newly established educational platforms, traditional KT models often fail due to the scarcity of student interaction data. To address this challenge, we leverage interaction data from existing education systems. Yet, these data sources differ substantially in subjects, exercises, and knowledge concepts.
We formulate this problem as a multi-source domain adaptation scenario for cold-start knowledge tracing and propose a novel domain-generalizable framework, DGKT. Our approach constructs generalized concept prototypes to unify semantic representations across domains and reduces domain discrepancies through a Sequence Instance Normalization (SeqIN) strategy. To better capture temporal dependencies in student behavior, we further introduce a Relation-Aware Attention Encoder (RA-Encoder). We evaluate our method on five benchmark datasets with extremely limited data, where each target domain contains fewer than 300 labeled training samples. Experimental results show that DGKT framework achieves an average AUC improvement of 1.98% across domains. Furthermore, our proposed DGrKT model outperforms existing KT models with a notable 4.16% AUC improvement. The source code is publicly available at: https://anonymous.4open.science/r/DGKT-5330.
keywords:
@@
Knowledge tracing , intelligent education , multi-source domain adaptation , cold-start , knowledge concept
1 Introduction
@@
Over the past decades, the emergence of numerous online education systems has significantly transformed the educational landscape by offering remote learning environments and personalized guidance for users. These advancements have revolutionized the educational landscape, enabling platforms to cater to the individual learning needs of students more effectively [4, 21, 40]. In this context, knowledge tracing (KT) plays a crucial role as it allows education systems to monitor and evaluate the evolving knowledge states of students. Knowledge tracing (KT) specifically aims to accurately track a student’s knowledge progression, typically through the prediction of future academic performance based on their historical interactions [2, 6, 39]. This enables online education platforms to better assess students’ comprehension and provide tailored assistance for learners [28, 30].
@@
Figure 1: An intuitive explanation for DGKT. Existing systems contain abundant interaction records while new education system contains only a few interactions. Our goal is to first train a versatile knowledge tracing model to acquire meta-knowledge from existing education systems and then apply it to the new system.
A number of KT models have demonstrated their effectiveness [10, 15, 1, 19, 26, 32]. However, these models typically require a large volume of student interactions to sufficiently train a KT model for a particular subject area. In the majority of knowledge tracing datasets, students’ problem-solving records typically have a scale of several hundred thousand or more. For example, there are 330, 116 interactions in ASSISTment 2009 to train a KT model sufficiently while it is challenging for newly developed education systems to accumulate such a substantial volume of data at the outset. In practical scenarios, the development of a new online education system or the introduction of a new question bank within an existing platform often lacks the extensive data on student interactions needed for effective KT model training [23, 35]. This scarcity of interactions results in the challenge of insufficient training data. As illustrated in Fig. 2, common KT models suffer a notable AUC degradation when confronted with limited data sizes.
Figure 2: The AUC result for three existing KT models (DKT [28], AKT [15], SANIT [10]) on ASSISTment 2009 with different data sizes where the proportion of available data gradually decreases.
@@
In cases where student interaction records are scarce in the new system, numerous student interaction records from various online education platforms provide valuable insights. These records may originate from different sources, yet they mirror the process of knowledge acquisition. Extracting overarching cognitive patterns from these varied student interaction sequences is beneficial for knowledge tracing in new systems. However, interaction sequences from different education systems differ a lot, so it is unfeasible to directly use them as training data.
Thus, we innovatively formulate the dilemma of the new education system as a Multi-source Domain Adaptation setting for cold-start KT, as shown in Fig. 1. Sufficient student interaction records from various online education domains, such as Algebra and Geometry, are treated as source domains. Meanwhile, the limited interactions from the new system, such as those in Probability, are treated as the target domain. Our goal is to train a domain-generalizable model that extracts meta-knowledge from the diverse student interactions across source domains, allowing for a seamless transfer of this knowledge to the target domain..
Figure 3: Kernel Density Estimate (KDE) for concept embedding from different education systems, i.e., ASSIST15, ALGEBRA05, and Junyi.
@@
To realize this goal, we propose a novel Domain Generalizable Knowledge Tracing (DGKT) framework. Notably, DGKT faces two primary challenges: 1) significant discrepancy among the source domains and 2) scarce student interactions within the target domain.
For the problem of significant distribution discrepancy among the source domains, as clearly illustrated in Fig. 3, we perform a kernel density estimation for the knowledge concepts from various domains, which reveals significantly distinct distributions for the concept embeddings from different domains. To overcome this challenge, we propose the concept aggregation—an algorithm that aggregates concept embeddings from different domains into concept prototypes for subsequent analysis of students’ interaction sequences.
@@
For the problem of scarce student interactions within the target domain, we recognize the challenge of learning accurate embeddings of the target question with limited interaction data. Consequently, we design a unique concept representation for the target domain that can effectively adapt to target domain with a few training data. Moreover, we propose a Relation-Aware Attention Encoder (RA-Encoder) that fully leverage the relation of the exercises in target domain. The contributions of this paper are summarized as follows:
  • 1.
    A novel domain-generalizable framework for cold-start KT. This issue is valuable but insufficiently studied. We formulate the multi-source domain adaptation setting and propose the DGKT framework that can quickly adapt to the target domain with minimal data. The DGKT framework can be integrated into existing KT models.
  • 2.
@@
    Innovative concept aggregation for reducing domain discrepancy. This addresses the significant challenge posed by the vast differences in concepts across various domains. Through concept aggregation, similar concepts are aggregated into a concept prototype.
  • 3.
@@
    The Relation-Aware Knowledge Encoder with better generalizability. We propose a relation-aware knowledge encoder (RA-Encoder) which fully leverages relational information within the target domain. By integrating RA-Encoder into DGKT framework, we construct the Domain-Generalizable Relation-aware Knowledge Tracing model (DGrKT). Our specially designed model demonstrates superior performance in multi-source domain adaptation for cold-start knowledge tracing.
@@
After conducting extensive experiments on five benchmark datasets, our DGKT framework achieves an average AUC improvement of 2.13%. Moreover, our DGrKT demonstrates the best performance, with an AUC improvement of 4.16% over five existing KT methods on multi-source domain adaptation for cold-start knowledge tracing tasks.
2 Related Work
Here, we introduce the related work including knowledge tracing, cold start and multi-source domain adaptation.
2.1 Knowledge Tracing
@@
Recently, many deep models for KT have emerged. The earliest deep model for KT is Deep Knowledge Tracing (DKT) [28]. DKT applies recurrent neural networks (RNNs) and outperforms traditional KT models. Many variants of DKT are proposed afterwards such as DKT+ [38]. Exercise-Enhanced Recurrent Neural Network (EERNN) [31] is proposed for student performance prediction by taking full advantage of both student exercising records and the text of each exercise. Another typical deep KT model is Separated Self-AttentIve Neural Knowledge Tracing (SAINT) [10] which applies Transformer to KT tasks. Attentive Knowledge Tracing (AKT) [15] is another attention-based KT model using a novel monotonic attention mechanism that relates a learner’s future responses to assessment questions to their past responses. Learning Process-consistent Knowledge Tracing (LPKT) [29] monitors students’ knowledge state through directly modeling their learning process. DASKT [32] incorporates simulated affective states into KT, improving interpretability and performance. LefoKT [5] introduces relative forgetting attention to decouple forgetting from item relevance and improve attention-based KT models’ ability to handle ever-growing interaction sequences. AdaptKT [9], delved into domain adaptation for knowledge tracing with text information and relatively sufficient data. As far as we know, few works focus on the data scarcity issue in online education systems, and we innovatively attempt to train a generalizable KT model.
Currently, few studies have delved into the data scarcity issue in knowledge tracing. Some prior efforts have aimed to address the sparse problem that students tend to interact with only a small set of questions, using approaches like pre-trained question embeddings [24] or contrastive learning [30, 37, 36]. In cases where interactions are severely limited, these approaches might not be effective in learning question embeddings and students’ knowledge states, as they heavily rely on a larger amount of student interactions. The scarcity issue of student interactions in new educational systems presents a significant challenge for knowledge tracing.
2.2 Cold Start and Multi-Source Domain Adaptation
The cold-start problem, where data is scarce, is a significant challenge in various real-world applications such as recommender systems and personalized learning. One promising solution to this issue is source-free domain adaptation (SFDA), which leverages knowledge from source domains with abundant data to assist target domains with limited data. SHOT [3] introduces a novel method to address this problem by employing a clustering-based pseudo-labeling technique and incorporating information maximization loss to align target domain features with the pretrained source model. In a similar vein, [20] highlight the trade-offs between discriminability and transferability and propose a strategy that combines original and translated samples using mix-up to enhance model performance.
When data from multiple source domains is available, Multi-source Domain Adaptation (MSDA) becomes a more effective approach, enabling knowledge transfer from various domains. In this context, [3] expands the SFDA framework by combining the outputs from different source models with learnable weights, and adapting the models using weighted information maximization. CAiDA [12] builds upon this idea by integrating a pseudo-label generator based on confident-anchor induction, further improving the adaptation process. DATE [17] takes a different approach by evaluating the transferability of source models through a Bayesian framework, quantifying the similarity between domains via a multi-layer perceptron. Bi-ATEN [22] introduces a tuning-free bi-level attention ensemble to adapt multiple source models to an unlabeled target domain without source data.
3 Domain-Generalizable Knowledge Tracing
@@
In this section, we elaborate on the proposed DGKT framework. Firstly, we formulate the KT tasks and their multi-source domain adaptation setting. Subsequently, we present the DGKT architecture. Furthermore, we introduce the process of concept aggregation. Finally, we introduce the RA-Encoder, which is specifically designed to address the multi-source domain adaptation problem.
3.1 Problem Definition
@@
A KT task with few sequences. In a KT task, there are very few sequences of interactions for training, denoted as $I=\{(q_{1},r_{1}),(q_{2},r_{2}),...,(q_{T},r_{T})\}$. Here, $T$ represents the length of the sequence, $q_{t}\in\mathbb{N}^{+}$ corresponds to the question ID of the $t$-th interaction ($t\leq T$), and $r_{t}\in\{0,1\}$ means the correctness of the student’s answer to the question $q_{t}$. These questions are associated with $n_{c}$ concepts that students need to master. The objective is to train a KT model capable of mining the knowledge state of students and predicting the probability that a student will answer the next question correctly, denoted as $P(r_{n+1}|q_{n+1},I)$. Since data scarcity affects model training, we resort to other available sequences as auxiliary source domain data for multi-source domain adaptation.
@@
Multi-source domain adaptation setting for cold-start KT. In the context of multi-source domain adaptation for cold-start knowledge tracing, there are $N$ auxiliary source domains $\{D_{s}^{i}|1\leq i\leq N\}$ and a target domain $D_{t}$. Each source domain $D_{s}^{i}$ consists of $m_{i}$ sequences of student interactions, denoted as $D_{s}^{i}=\{I_{j}^{i}|1\leq j\leq m_{i}\}$ along with a set of questions $\{q_{1},q_{2},...,q_{n_{q_{i}}}\}$ and a set of concepts $\{c_{1},c_{2},...,c_{n_{c_{i}}}\}$, where $n_{q_{i}}$ represents the number of questions, $n_{c_{i}}$ represents the number of concepts. The target domain, on the other hand, contains only $m_{t}$ student interaction sequences, represented as $D_{t}=\{I_{j}^{t}|1\leq j\leq m_{t}\}$, with $m_{t}\ll m_{i}$. Importantly, the data $D_{t}$ from the target domain is unseen during the meta-training phase. The objective is to train a generalized KT model by leveraging all the source domain data $\{D_{s}^{i}|1\leq i\leq N\}$, and subsequently adapt this model to the target domain data $D_{t}$ for knowledge tracing.
@@
3.2 DGKT Architecture
The process outlined in Fig. 4 illustrates the key steps of our DGKT approach. Initially, we employ the feature embedding module to transform the interaction sequences into corresponding embedding sequences, followed by the knowledge state encoder that generates the hidden knowledge states. Then, the knowledge states are decoded by the knowledge state decoder, finally producing the anticipated probability of a student providing a correct response to the subsequent question.
Feature embedding module. In feature embedding module, our objective is to convert the sequence of questions and responses into the embedding sequence. For each domain, we have a concept matrix $\textbf{Q}\in\mathbb{R}^{n_{c}\times n_{q}}$, where $n_{q}$ represents the total number of question IDs and $n_{c}$ represents the total number of knowledge concept IDs. The element $\textbf{Q}_{ij}$ is set to 1 if the $j$-th question is related to the $i$-th knowledge concept. Then, we learn the embedding $e_{q_{t}}\in\mathbb{R}^{d}$ of question $q_{t}$ ($d$ is the dimension of the embedding), by averaging the embeddings $e_{i}$ of all concepts associated with the question, which can be written by
$e_{q_{t}}=\frac{\sum_{i}\mathbb{I}(\textbf{Q}_{iq_{t}}=1)e_{i}}{\sum_{i}\mathbb{I}(\textbf{Q}_{iq_{t}}=1)},$ (1)
where $\mathbb{I}(\cdot)$ represents the indicator function, and $e_{i}\in\mathbb{R}^{d}$ is a learnable vector representing the $i$-th concept.
Afterward, we leverage the response $r_{t}$ of the question to construct the question-response embedding $e_{qr_{t}}\in\mathbb{R}^{2d}$ as follows: